# This Version: March 9, 2024

### v1

@@ 2 General Framework
We begin by framing the environment for a Markov Decision Process (MDP) with model misspecification. While the conceptual framework is patterned after the finite (states and actions) environment of EP, we adopt the setting of ADGK since it applies to infinite settings; an environment that naturally features many important applications in economics. After outlining the contents of a misspecified MDP and the relevant Berk-Nash equilibrium concept, we provide an overview of the order-theoretic methods that are instrumental for our results. A refresher for these concepts and methods is provided in Appendix A for the reader’s convenience.
At the start of each period $t=0,1,2,\ldots,$ the agent observes a state realization, $s_{t}\in\mathbb{S},$ and then chooses an action, $x_{t}\in\mathbb{X}.$ Given a transition probability function $Q(\cdot|s_{t},x_{t}),$ the state and action together determine the distribution of the next period state, $s_{t+1}.$ The per-period payoff function is a mapping $u:\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}.$ The agent maximizes expected discounted utility (discount factor $10<\beta<$ 1) by choosing a feasible sequence of policy functions $\{x_{t}\}_{t=1}^{\infty}$ that solves the following problem,
$V(s_{0})=\max_{\{x_{t}\}_{t=0}^{\infty}}\mathbb{E}_{Q}\Bigg{[}\sum_{t=0}^{\infty}\beta^{t}u(s_{t},x_{t})\Bigg{]},\ t=0,1,2,\ldots.$ (2)
@@
$V(s)=\max_{x\in\mathbb{X}}\Bigl{\{}u\left(s,x\right)+\beta\int_{\mathbb{S}}V\left(s^{\prime}\right)Q\left(ds^{\prime}\mid s,x\right)\Bigl{\}},$ (3)
@@
$G(s)\equiv\Bigl{\{}\hat{x}\in\mathbb{X}|\\ \ \hat{x}\in\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\displaystyle\Bigl{\{}u(s,x)+\beta\int_{S}V(s^{\prime})Q(\mathrm{d}s^{\prime}|s,x)\Bigl{\}}\Bigl{\}},$ (4)
where action $\hat{x}$ is optimal given state $s$. We summarize this environment in the following definition of a MDP. [^16] The framework can be extended with minor modifications to allow for the dependence of feasible set of actions on the state variable; see the Esponda and Pouzo (2015) working paper for details.
@@
A subjective Markov Decision Process (herafter, SMDP) is an MDP, $\left\langle\mathbb{S},\mathbb{X},q_{0},Q\right.$, $u\rangle$, and a non-empty family of transition probability functions, $\mathcal{Q}_{\Theta}=\left\{Q_{\theta}:\theta\in\Theta\right\}$, where each transition probability function $Q_{\theta}:\mathbb{S}\times\mathbb{X}\rightarrow\mathcal{M}_{1}(\mathbb{S})$ is indexed by a parameter value $\theta\in\Theta\subseteq\mathbb{R}$. A SMDP is said to be misspecified if $Q\notin\mathcal{Q}_{\Theta}.$
Notice that under this definition, an SMDP could be misspecified in several ways. For instance, the true transition function and the set of model transition functions may pertain to dissimilar families of probability distributions. Furthermore, even if the true model and set of models belong to the same family of distributions, they can be misspecified if the support, the range of possible values of the parameter, of the models is different from that of the true distribution. [^18] See Examples 1 and 2 for an instance for the former, and Example 3 for the latter. Den Haan and Drechsel (2019) explore more on forms of model misspecification and their relevance to econometric methods with macroeconomic models. Our next definition requires the primitives of the SMDP to satisfy certain regularity conditions.
@@
    (Continuity) The mappings $(s,x)\to Q(\cdot|s,x)$ and $(\theta,s,x)\to Q_{\theta}(\cdot|s,x)$ are continuous in the Prokhorov metric.
@@
Condition (i) is a standard technical condition and requires the transition functions, both true and subjective, to be continuous. Condition (ii) requires that there always exists some model $Q_{\theta}$ that accounts for every observation from the true distribution, $Q.$ Condition (iii) places an uniform integrability requirement to deal with distributions with infinite support. Lastly, condition (iv) requires the parameter space $\Theta$ to be compact. This is an essential requirement for the existence of best-fit models as will be evident in the next definition. We next measure the cost of misspecification in terms of the well-known measure of relative entropy, the Kullback-Liebler divergence.
@@
For a given true transition function $Q$ and model transition function $Q_{\theta},$ the Kullback-Liebler divergence, $\mathcal{D}_{\mathrm{KL}},$ of $Q_{\theta}$ with respect to $Q$ is defined as,
$\mathcal{D}_{\mathrm{KL}}\big{(}Q(s,x),Q_{\theta}(s,x)\big{)}=\mathbb{E}_{Q(\cdot|s,x)}\left[\ln\big{(}D_{\theta}(s^{\prime}|s,x)\big{)}\right],$ (5)
where $\big{(}D_{\theta}(s^{\prime}|s,x)\big{)}$ is the Radon-Nikodym derivative of $Q$ with respect to $Q_{\theta},\ \theta\in\Theta.$ Then for any distribution over states and actions, $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}),$ and model parameter, $\theta\in\Theta$, the weighted Kullback-Leibler divergence is a mapping $K_{Q}:\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\Theta\to\bar{\mathbb{R}}_{\geq 0}$ such that
@@
However, if the SMDP is misspecified, that is, if $Q$ is not contained within the set $\mathcal{Q}_{\Theta}$, then in Definition (5), condition (a) asserts that action $x$ is optimal in the MDP$(\bar{Q}_{\mu^{*}})$ when the equilibrium transition function is a weighted combination of model transition functions, with weights given by $\mu^{*}$. Condition (b) mandates that these beliefs $\mu^{*}$ are determined by minimizing the weighted relative entropy between the true transition function and the model transition functions, where the weights are the Berk-Nash equilibrium, $m^{*}$. Lastly, condition (c) stipulates that the marginal distributions over states, $m^{*}_{\mathbb{S}}$ is an invariant distribution. Notice that in condition (c), the invariant distribution is dependent on the optimal actions $x$ via the true transition function, $Q$. EP show that under moderate conditions, the learning illustration with misspecified models sketched in Equation (1) converges to the Berk-Nash equilibrium. [^22] For finite state and action spaces, EP shows that convergence happens under a mild identification assumption. For infinite (but compact) state and action spaces, ADGK demonstrate that convergence occurs under identification only under the total-variation norm on the space of state and action distributions. The equilibrium serves as a prediction of the steady-state behavior in where $m^{*}$ is the steady-state distribution over states and actions, and $\mu^{*}$ is the limit of the sequence of posteriors of a Bayesian learner with misspecified models.
@@
$V(s,\mu^{*},p)=\max_{x\in\mathbb{X}}\Bigl{\{}u(s,x)+\beta\int_{S}V(s^{\prime},\mu^{*},p)\bar{Q}_{\mu^{*}}(ds^{\prime}|s,x),\Bigl{\}}\mbox{ and }$ (7)
@@
$G(s,\mu^{*},p)\equiv\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\Bigl{\{}u(s,x)+\beta\int_{S}V(s^{\prime},\mu^{*},p)\bar{Q}_{\mu^{*}}(ds^{\prime}|s,x)\Bigl{\}}.$ (8)
Therefore, in line with Definition 5, the set of Berk-Nash equilibrium for a given primitive $p$ is given by the set of fixed points of the equilibrium mapping $T$,
@@
where $T:Z\times P\twoheadrightarrow 2^{Z}$ is a set-valued function on the space of probability measures on states and actions and the set of parameters, $Z=\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\mathcal{M}_{1}(\Theta)$ and $P=<u,\beta,Q,Q_{\Theta},\Theta>$ are our primitives. [^23] For the ease of the reader, the full expression for $T$ is defined in the Proofs section, p.30.
For inquiring the comparative statics behavior of the equilibrium objects in $\Lambda(p)$ with respect to the primitives of the environment, we need to define corresponding orders for the equilibrium objects and the primitives. For vectors $x$ and $y$ in $\mathbb{R}^{n}$, we use the following convention: “$x\geq y$” means $x_{i}\geq y_{i}$ in every component, “$x>y$” means $x\geq y$ and $x\neq y$, and “$x\gg y$” means $x_{i}>y_{i}$ in every component. A real-valued function, $f:X\rightarrow\mathbb{R},$ is said to be increasing if for $x\geq y$ in the component-wise order, $f(x)\geq f(y),$ and convex if its domain is a convex set and for all $x,y$ in its domain, and all $\lambda\in[0,1]$, we have $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y).$ [^24] The interested reader is referred to Shaked and Shanthikumar (2007) for more details concerning univariate and multivariate orders. The results in this paper could potentially be extended to admit various orders. We follow Shaked and Shanthikumar (2007) in ranking probability measures on the state and action space, and the parameter space.
@@
Let $\mathcal{M}(X)$ denote the space of probability measures defined on a compact subset of $X\subseteq\mathbb{R}^{n}$. [^25] To be clear, $X$ denotes a general set, not particularly the set of actions, $\mathbb{X}.$ For any two measures, $\mu$ and $\nu$ in $\mathcal{M}(X),$
@@
Our setting fundamentally diverges from the setting in Acemoglu and Jensen (2015) in several key aspects. In their work, they examine an infinite-horizon dynamic economy featuring a continuum of agents subject to idiosyncratic shocks, whereas our setting deals with a single agent facing an infinite-horizon dynamic problem in which shocks are influenced by action choices. Most importantly, our analysis intertwines the agent’s decision-making process with learning under model misspecification, whereas in Acemoglu and Jensen (2015), the agents possess knowledge of the accurate distribution of idiosyncratic shocks, thus obviating the need for any learning component. This also furthers tailoring their techniques to our framework.
Towards this end, we shall show (Lemmas 1 and 2 in Appendix B) that our equilibrium mapping $T$ is monotonic in the sense defined below.
@@ 3 Motivating Examples
We now present three examples of misspecified environments that illustrate the framework of this paper. Our first example is that of a misspecified AR(1) inference process where actions have no role. It is purely an inference problem. [^30] Per se, there is no decision process involved here. One could think of an economic situation where the flow of utility is constant over time, and therefore, actions play no role in either inference, nor, payoffs. Examples 2 and 3 are more substantive in that the choice of actions plays a role in learning, and vice-versa. They are based on EP and cover effort provision problems with misspecified effort-task dynamics and consumption-savings problem with misspecified wealth process, respectively. While Example 2 has an explicit analytical solution and is, therefore, used to outline the mechanics for comparative statics properties of a Berk-Nash equilibrium, the solution for Example 3 is intractable and, therefore, the most substantive for the applicability of this paper’s techniques. [^31] We omit the full analysis of Examples 2 and 3 since they are already covered in the aforementioned papers.
@@
$\theta^{*}=\displaystyle\int_{\mathbb{S}}\hat{\theta}(s)m^{*}_{\mathbb{S}}=\rho+\displaystyle\int_{\mathbb{S}}\dfrac{(\mu_{1}+\mu_{2})}{s}m^{*}_{\mathbb{S}}.$
Now, let $\mu_{1}+\mu_{2}=0.$ That is, the mean of the components of the true distribution cancel each other. In this case, a Bayesian agent, starting from any prior on their models $Q_{\Theta}$, would eventually converge in the Berk-Nash equilibrium to the true persistence parameter $\rho$ and yet continue to remain misspecified. For this case, the comparative statics of the steady state inferred $\theta^{*}$ with respect to the AR(1) parameter of the true process is one-to-one in a trivial way. [^33] For the case when $\mu_{1}+\mu_{2}\neq 0,$ the expression for $\theta^{*}$ is potentially intractable since the expectation with respect to the Berk-Nash equilibrium $m^{*}$ may simply not exist. The inspiration for the mixture normal distribution comes from a paper by Csaba and Szoke (2023) who look at learning when the set of likelihoods is misspecified.
@@
In this setting of a dynamic effort provision problem, we follow EP in characterizing the Berk-Nash equilibrium. The equilibrium for this instance is a probabilistic prediction for the agent’s long run outcome to be either a success or a failure, along with their inference of their ability. This example is related to the growing literature on effort provision, as exemplified in Deimen and Wirtz (2022).
@@
For this case, the Berk-Nash equilibrium is characterized by an optimal policy function, $x^{m}=A_{z}\left(\beta^{m}\right)y=\displaystyle\dfrac{0.5\delta\beta^{m}}{\left(1-\delta\beta^{m}\right)z+0.5\delta\beta^{m}}y$, where there exists a corresponding inferred model $\beta^{m}\in\left(0,\beta^{*}\right)$. That is, the misspecified agent underestimates the return on her saving, $\beta^{m}<\beta^{*},$ and therefore, undersaves in the Berk-Nash equilibrium, $x^{m}<x^{*}.$ [^35] The intuition behind this is that a higher preference shock $z$ is associated with a lower saving proportion, $A_{z}(\beta^{m}),$ since it leads to a lower inferred belief of return on saving, $\beta^{m},$ since the preference and productivity shocks are positively correlated. Notice however that there is no closed-form expression for $\beta^{m},$ that outlines its dependence on the environment primitives, and therefore, how the inferred return, the equilibrium policy function, and the invariant distribution respond to the primitives of the environment is an open question. [^36] From EP: The inferred model at the stationary Berk-Nash equilibrium solves the following equation, $\hat{\beta}(A_{z}(\beta))=\beta^{*}+\gamma^{*}\frac{\operatorname{Cov}\left(z,\ln A_{z}(\beta)\right)}{\operatorname{Var}\left(\ln A_{z}(\beta)\right)+\operatorname{Var}(\ln Y)},$ (13) where the covariance and variance are taken with respect to the true distribution, $Q(y^{\prime},z^{\prime}|y,z).$ Under the parametric assumptions of $\gamma^{*}>0,$ and $Cov(z,A_{z})<0$, EP establish that there exists a $\beta^{m},$ the inferred model in the Berk-Nash equilibrium. Further, in EP, notice that the inferred model depends on the proportion, $A_{z},$ instead of the distribution over states and actions, which is unlike the case required as per the conditions of Berk-Nash equilibria.

### v2

@@ 2 General Framework
We begin by framing the environment for a Markov Decision Process (MDP) with model misspecification. While the conceptual framework is patterned after the finite (states and actions) environment of EP, our setting is based on Euclidean spaces since it allows infinite settings that naturally feature many important economic applications. After outlining the contents of a misspecified MDP and the relevant Berk-Nash equilibrium concept, we provide an overview of the order-theoretic methods that are instrumental for our results. A refresher for these concepts and methods is provided in Appendix A for the reader’s convenience.
At the start of each period $t=0,1,2,\ldots,$ the agent observes a state realization, $s_{t}\in\mathbb{S},$ and then chooses an action, $x_{t}\in\mathbb{X}.$ Given a transition probability function $Q(\cdot|s_{t},x_{t}),$ the state and action together determine the distribution of the next period state, $s_{t+1}.$ The per-period payoff function is a mapping $u:\mathbb{S}\times\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}.$ The agent maximizes expected discounted utility (discount factor $10<\beta<$ 1) by choosing a feasible sequence of policy functions $\{x_{t}\}_{t=1}^{\infty}$ that solves the following problem,
$V(s_{0})=\max_{\{x_{t}\}_{t=0}^{\infty}}\mathbb{E}_{Q}\Bigg{[}\sum_{t=0}^{\infty}\beta^{t}u(s_{t},x_{t},s_{t+1})\Bigg{]},\ t=0,1,2,\ldots.$ (2)
@@
$V(s)=\max_{x\in\mathbb{X}}\Bigl{\{}\int_{\mathbb{S}}\{u\left(s,x,s^{\prime}\right)+\beta V\left(s^{\prime}\right)\}Q\left(ds^{\prime}\mid s,x\right)\Bigl{\}},$ (3)
@@
$G(s)\equiv\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\displaystyle\Bigl{\{}\int_{\mathbb{S}}\{u\left(s,x,s^{\prime}\right)+\beta V(s^{\prime})\}Q(\mathrm{d}s^{\prime}|s,x)\Bigl{\}}.$ (4)
We summarize this environment in the following definition of a MDP. [^18] The framework can be extended with minor modifications to allow for the dependence of feasible set of actions on the state variable; see the Esponda and Pouzo (2015) working paper for details.
@@
A Subjective Markov Decision Process (hereafter, SMDP) is an MDP, $\left\langle\mathbb{S},\mathbb{X},q_{0},Q\right.$, $u\rangle$, and a non-empty family of transition probability functions, $\mathcal{Q}_{\Theta}=\left\{Q_{\theta}:\theta\in\Theta\right\}$, where each transition probability function $Q_{\theta}:\mathbb{S}\times\mathbb{X}\rightarrow\mathcal{M}_{1}(\mathbb{S})$ is indexed by a parameter value $\theta\in\Theta\subseteq\mathbb{R}$ [^20] Our focus here is on an uni-dimensional parameter space $\mathbb{R}$ to align with applications commonly found in the literature, however, the results in this paper are applicable to $\mathbb{R}^{d},$ with slight modifications. The reader is also referred to the discussion on multi-dimensional parameter spaces in Section 7. A SMDP is said to be misspecified if $Q\notin\mathcal{Q}_{\Theta}.$
Notice that under this definition, an SMDP could be misspecified in several ways. For instance, the true transition function and the set of model transition functions may pertain to dissimilar families of probability distributions. Alternatively, even if the true model and set of models belong to the same family of distributions, they can be misspecified if the support, the range of possible values of the parameter, of the models is different from that of the true distribution. [^21] See Examples 1 and 2 for an instance for the former, and Example 3 for the latter. Den Haan and Drechsel (2019) explore more on forms of model misspecification and their relevance to econometric methods with macroeconomic models. Our next definition requires the primitives of the SMDP to satisfy certain regularity conditions.
@@
    (Continuity) The mappings $(s,x)\to Q(\cdot\mid s,x)$ and $(\theta,s,x)\to Q_{\theta}(\cdot\mid s,x)$ are continuous in the Prokhorov metric, and the density function $D_{\theta}\left(s^{\prime}\mid s,x\right)$ is jointly continuous on the set
    $\left\{\left(\theta,s^{\prime},s,x\right):Q(s,x)\text{ is dominated by }Q_{\theta}(s,x)\right\},$
    where $\big{(}D_{\theta}(s^{\prime}\mid s,x)\big{)}$ is the Radon-Nikodym derivative of $Q$ with respect to $Q_{\theta},\ \theta\in\Theta.$ [^22] The Prokhorov metric is used to measure the closeness of probability measures based on their tightness and weak convergence. For two transition functions $Q_{1}$ and $Q_{2}$, the Prokhorov metric is defined as: $d(Q_{1},Q_{2})=\inf\{\epsilon>0:Q_{1}(A)\leq Q_{2}(A^{\epsilon})+\epsilon\text{ and }Q_{2}(A)\leq Q_{1}(A^{\epsilon})+\epsilon,\ \forall A\in\mathcal{B}(\mathbb{X})\},$ where $A^{\epsilon}$ is the $\epsilon$-neighborhood of $A$, given by: $A^{\epsilon}=\{x\in\mathbb{X}:d(x,a)<\epsilon\text{ for some }a\in A\}.$ This metric is particularly well-suited for compact state spaces.
@@
Condition (i) is a standard technical condition and requires the transition functions, both true and subjective, to be continuous. Condition (ii) requires that there always exists some model $Q_{\theta}$ that accounts for every observation from the true distribution, $Q.$ Condition (iii) places an uniform integrability requirement to deal with distributions with infinite support. Lastly, condition (iv) requires the parameter space $\Theta$ to be compact. This is an essential requirement for the existence of best-fit models as will be evident in the next definition. We next measure the extent of misspecification in terms of the well-known measure of relative entropy, the Kullback-Leibler divergence.
@@
For a given true transition function $Q$ and model transition function $Q_{\theta},$ the Kullback-Leibler divergence, $\mathcal{D}_{\mathrm{KL}},$ of $Q_{\theta}$ with respect to $Q$ is defined as,
$\mathcal{D}_{\mathrm{KL}}\big{(}Q(s,x),Q_{\theta}(s,x)\big{)}=\mathbb{E}_{Q(\cdot|s,x)}\left[\ln\big{(}D_{\theta}(s^{\prime}|s,x)\big{)}\right].$ (5)
Then for any distribution over states and actions, $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}),$ and model parameter, $\theta\in\Theta$, the weighted Kullback-Leibler divergence is a mapping $K_{Q}:\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\Theta\to\bar{\mathbb{R}}_{\geq 0}$ such that
@@
However, if the SMDP is misspecified, that is, if $Q$ is not contained within the set $\mathcal{Q}_{\Theta}$, then in Definition (5), condition (a) asserts that action $x$ is optimal in the MDP$(\bar{Q}_{\mu^{*}})$ when the equilibrium transition function is a weighted combination of model transition functions, with weights given by $\mu^{*}$. Condition (b) mandates that these beliefs $\mu^{*}$ are determined by minimizing the weighted relative entropy between the true transition function and the model transition functions, where the weights are the Berk-Nash equilibrium, $m^{*}$. Lastly, condition (c) stipulates that the marginal distributions over states, $m^{*}_{\mathbb{S}}$ is an invariant distribution. Notice that in condition (c), the invariant distribution is dependent on the optimal actions $x$ via the true transition function, $Q$. EP show that under moderate conditions, the learning illustration with misspecified models sketched in Equation (1) converges to the Berk-Nash equilibrium. [^26] EP, in their paper, demonstrate the usefulness of their equilibrium concept by providing a learning foundation for finite spaces (Theorems 2 and 3), under either positive visitation or identification. The identification condition also holds in our setting, courtesy of Assumption 3. Anderson, Duanmu, Ghosh, and Khan (2022) have partially addressed the convergence problem for infinite spaces in an ArXiv version (v2) of their manuscript, available here: [https://arxiv.org/pdf/2206.08437v2; see Section C.2 in Appendix, pages 48–56]. For infinite (but compact) state and action spaces, they demonstrate that convergence holds under identification, but only in the total-variation norm on the space of state and action distributions. The equilibrium serves as a prediction of the steady-state behavior in where $m^{*}$ is the steady-state distribution over states and actions, and $\mu^{*}$ is the limit of the sequence of posteriors of a Bayesian learner with misspecified models.
@@
$V(s,\mu^{*},p)=\max_{x\in\mathbb{X}}\Bigl{\{}\int_{\mathbb{S}}\{u\left(s,x,s^{\prime}\right)+\beta V(s^{\prime},\mu^{*},p)\}\bar{Q}_{\mu^{*}}(ds^{\prime}|s,x),\Bigl{\}}\mbox{ and }$ (7)
@@
$G(s,\mu^{*},p)\equiv\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\Bigl{\{}\int_{\mathbb{S}}\{u\left(s,x,s^{\prime}\right)+\beta V(s^{\prime},\mu^{*},p)\}\bar{Q}_{\mu^{*}}(ds^{\prime}|s,x)\Bigl{\}}.$ (8)
Therefore, in line with Definition 5, the set of Berk-Nash equilibriua for a given primitive $p$ is given by the set of fixed points of the equilibrium mapping $T$,
@@
where $T:Z\times P\twoheadrightarrow 2^{Z}$ is a set-valued function on the space of probability measures on states and actions and the set of parameters, $Z=\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\mathcal{M}_{1}(\Theta)$ and $P=<u,\beta,Q,Q_{\Theta},\Theta>$ are our primitives.
For inquiring the comparative statics behavior of the equilibrium objects in $\Lambda(p)$ with respect to the primitives of the environment, we need to define corresponding orders for the equilibrium objects and the primitives. For vectors $x$ and $y$ in $\mathbb{R}^{n}$, we use the following convention: “$x\geq y$” means $x_{i}\geq y_{i}$ in every component, “$x>y$” means $x\geq y$ and $x\neq y$, and “$x\gg y$” means $x_{i}>y_{i}$ in every component. A real-valued function, $f:X\rightarrow\mathbb{R},$ is said to be increasing if for $x\geq y$ in the component-wise order, $f(x)\geq f(y),$ and convex if its domain is a convex set and for all $x,y$ in its domain, and all $\lambda\in[0,1]$, we have $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y).$ [^27] The interested reader is referred to Shaked and Shanthikumar (2007) for more details concerning univariate and multivariate orders. We follow Shaked and Shanthikumar (2007) in ranking probability measures on the state and action space, and the parameter space.
@@
Let $\mathcal{M}(X)$ denote the space of probability measures on a compact subset $X\subseteq\mathbb{R}^{k}$ for some finite $k\in\mathbb{N}$. For any two measures, $\mu$ and $\nu$ in $\mathcal{M}(X),$
@@
Our setting differs from that of Acemoglu and Jensen (2015) in several key aspects. Their work examines the steady-state outcomes of an infinite-horizon dynamic economy with a continuum of agents experiencing idiosyncratic (exogenous) shocks, whereas we focus on the steady-state outcomes of a single agent facing an infinite-horizon dynamic problem, where shocks depend on the agent’s action choices and are therefore endogenous. Furthermore, our setting features a Bayesian agent learning with misspecified models of the shock process, unlike Acemoglu and Jensen (2015), which assumes that agents know the true distribution of idiosyncratic shocks, thereby removing the need for any learning component. As a result, their techniques require adaptations to suit our framework (see Auxiliary Lemmas 1 and 2).
Towards this end, we shall show (Auxiliary Lemmas 1 and 2 in Appendix B) that our equilibrium mapping $T$ is monotonic in the sense defined below.
@@ 3 Motivating Examples
We now present three examples of misspecified environments that illustrate the framework of this paper. Our first example is that of a misspecified AR(1) inference process where actions have no role. It is purely an inference problem. [^32] Per se, there is no decision process involved here. One could think of an economic situation where the flow of utility is constant over time, and therefore, actions play no role in either inference, nor, payoffs. Examples 2 and 3 are more substantive in that the choice of actions plays a role in learning, and vice-versa. They are based on EP and cover effort provision problems with misspecified effort-task dynamics and consumption-savings problem with misspecified wealth process, respectively. While Example 2 has an explicit analytical solution and is, therefore, used to outline the mechanics for comparative statics properties of a Berk-Nash equilibrium, the solution for Example 3 is intractable and, therefore, the most substantive for the applicability of this paper’s techniques. [^33] We omit the full analysis of Examples 2 and 3 since they are already covered in the aforementioned papers (EP and ADGK).
@@
$\theta^{*}=\displaystyle\int_{\mathbb{S}}\hat{\theta}(s)m^{*}_{\mathbb{S}}=\rho+\displaystyle\int_{\mathbb{S}}\dfrac{(\mu_{1}+\mu_{2})}{2s}m^{*}_{\mathbb{S}}.$
Now, suppose $\mu_{1}+\mu_{2}=0.$ That is, the mean of the components of the true distribution cancel each other. In this case, a Bayesian agent, starting from any prior on their models $Q_{\Theta}$, would eventually converge in the Berk-Nash equilibrium to the true persistence parameter $\rho$ and yet continue to remain misspecified. For this case, the comparative statics of the steady state inferred $\theta^{*}$ with respect to the AR(1) parameter of the true process is one-to-one in a trivial way. [^35] For the case when $\mu_{1}+\mu_{2}\neq 0,$ the expression for $\theta^{*}$ is potentially intractable since the expectation with respect to the Berk-Nash equilibrium $m^{*}$ may simply not exist. The inspiration for the mixture normal distribution comes from a paper by Csaba and Szoke (2023) who look at learning when the set of likelihoods is misspecified.
@@
In this setting of a dynamic effort provision problem, we follow EP in characterizing the Berk-Nash equilibrium. The equilibrium for this instance is a probabilistic prediction for the agent’s long run outcome to be either a success or a failure, along with their inference of their ability. This example is related to the growing literature on effort provision, as modeled in Deimen and Wirtz (2022).
@@
For this case, the Berk-Nash equilibrium is characterized by an optimal policy function, $x^{m}=A_{z}\left(\beta^{m}\right)y=\displaystyle\dfrac{0.5\delta\beta^{m}}{\left(1-\delta\beta^{m}\right)z+0.5\delta\beta^{m}}y$, where there exists a corresponding inferred model $\beta^{m}\in\left(0,\beta^{*}\right)$. That is, the misspecified agent underestimates the return on her saving, $\beta^{m}<\beta^{*},$ and therefore, undersaves in the Berk-Nash equilibrium, $x^{m}<x^{*}.$ [^37] The intuition behind this is that a higher preference shock $z$ is associated with a lower saving proportion, $A_{z}(\beta^{m}),$ since it leads to a lower inferred belief of return on saving, $\beta^{m},$ since the preference and productivity shocks are positively correlated. Notice however that there is no closed-form expression for $\beta^{m}$ that outlines its dependence on the environment primitives, and therefore, how the inferred return, the equilibrium policy function, and the invariant distribution respond to the primitives of the environment is an open question. [^38] From EP: The inferred model at the stationary Berk-Nash equilibrium solves the following equation, $\hat{\beta}(A_{z}(\beta))=\beta^{*}+\gamma^{*}\frac{\operatorname{Cov}\left(z,\ln A_{z}(\beta)\right)}{\operatorname{Var}\left(\ln A_{z}(\beta)\right)+\operatorname{Var}(\ln Y)},$ (13) where the covariance and variance are taken with respect to the true distribution, $Q(y^{\prime},z^{\prime}|y,z).$ Under the parametric assumptions of $\gamma^{*}>0,$ and $Cov(z,A_{z})<0$, EP establish that there exists a $\beta^{m},$ the inferred model in the Berk-Nash equilibrium. Further, in EP, notice that the inferred model depends on the proportion, $A_{z},$ instead of the distribution over states and actions, which is unlike the case required as per the conditions of Berk-Nash equilibria.